OpenStudy (anonymous):

Use synthetic division to find the remainder when 4 - 5x - x^2 + 6x^3 is divided by 3x - 2 a) -1 b) 4 c) 0 d) 2

OpenStudy (anonymous):

I understand synthetic division, however I've never dealt with a problem where I have to divide by the x with a coefficient greater than 1.

OpenStudy (anonymous):

solve 3x-2=0 to find the number you are synthetically dividing by.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

x=2/3 ?

OpenStudy (anonymous):

Yep, then do synth.div. as normal (remember to arrange the coefficients in the correct order).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so do i change the 2/3 to a -2/3?

OpenStudy (anonymous):

No, just use the solution for x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If you were dividing by (x-2) then you would synthetically divided by 2, so for division by (3x-2) you synth.div. by 2/3.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so far, I have 4, -2.33, -2.55, and a remainder of 4.3

OpenStudy (amistre64):

you didnt put your poly in decreasing order of degrees

OpenStudy (amistre64):

also, consider factoring out the 3 form top and bottom ... might be useful

OpenStudy (anonymous):

thanks i'll try that out

OpenStudy (amistre64):

4 - 5x - x^2 + 6x^3 6/3 -1/3 -5/3 4/3 0 4/3 2/3 -2/3 ---------------------- 2/3 ) 2 1 -1 2/3 3(2/3) = 2 if im seeing it right

OpenStudy (anonymous):

yeh i got the same thing :D

OpenStudy (amistre64):

good, then we are either both right, or both wrong :)

OpenStudy (anonymous):

:3

OpenStudy (anonymous):